the perimeter of a rectangular pool is 140 yd. if the length is 20 yd less than twice the width, find the length and the width
Answer by mando271(42)

i'll use L for length and W for width.
you are given that L=2W-20.
and you need 2 lengths and 2 width to form a rectangle, so.
2W+2(2W-20)=140
2W+4W-40=140
6W=180
W=30
so width equals 30, now plug this in to the 1st equation to solve for L
L=2(30)-20
L=60-20
L=40